# LLM-Based Threat Detection and Prevention Framework for IoT Ecosystems

### v1

@@ Abstract
The increasing complexity and scale of the Internet of Things (IoT) have made security a critical concern. This paper presents a novel Large Language Model (LLM)-based framework for comprehensive threat detection and prevention in IoT environments. The system integrates lightweight LLMs fine-tuned on IoT-specific datasets (IoT-23, TON_IoT) for real-time anomaly detection and automated, context-aware mitigation strategies optimized for resource-constrained devices. A modular Docker-based deployment enables scalable and reproducible evaluation across diverse network conditions. Experimental results in simulated IoT environments demonstrate significant improvements over traditional security methods in detection accuracy, response latency, and resource efficiency. The proposed framework highlights the potential of LLM-driven, autonomous security solutions for future IoT ecosystems.
@@ I Introduction
The rapid growth of Internet of Things (IoT) ecosystems has introduced complex challenges in managing and optimizing system performance. IoT environments generate vast amounts of data, requiring efficient real-time processing and advanced operations such as resource allocation, predictive maintenance, and energy management. The diverse and dynamic nature of IoT devices further complicates these challenges, as devices often operate under heterogeneous conditions with varying computational and network capabilities [1]. The rapid expansion of IoT ecosystems introduces growing complexities in securing large-scale networks [2]. Traditional IoT security techniques often struggle to detect evolving and sophisticated threats. Recent advances in Large Learning Models (LLMs) have shown promise in improving security systems, offering adaptability to evolving threats, scalability across diverse IoT environments, and effectiveness in managing complex datasets. Recent studies demonstrate the growing potential of LLMs in cybersecurity applications. For example, the authors of [3] explored using domain-specific fine-tuning in LLMs for detecting zero-day vulnerabilities. Similarly, Patel et al. [4] proposed a hybrid approach combining LLMs with graph-based anomaly detection to secure IoT networks against emerging threats. The integration of LLMs with federated learning frameworks, as shown by Liu et al. [5], has also been successful in enhancing data privacy while maintaining robust security in distributed IoT environments. Moreover, Kumar et al. [6] highlighted the effectiveness of context-aware LLMs in identifying complex attack patterns specific to IoT ecosystems. The research by Smith et al. [7] underlined the importance of adaptive learning techniques for LLMs to ensure continuous improvement in threat detection and prevention. Our model presents an LLM-based security system tailored to IoT environments, focusing on a two-component approach for threat detection and prevention. The proposed system integrates the strengths of LLMs, utilizing prompt engineering, fine-tuning, and domain-specific training to identify and respond to security breaches in real-time. By leveraging IoT-specific datasets such as IoT-23 and TON_IoT, the system continuously adapts to new threats, ensuring it remains effective in dynamic environments. This framework represents a significant step toward overcoming the limitations of traditional solutions while addressing the unique challenges of integrating advanced AI models into IoT ecosystems. The main contributions of this work are the following:
@@ II Related Works
In recent years, IoT security has been an active research area, with numerous studies exploring Machine Learning (ML) techniques for anomaly detection and intrusion prevention. Several works have demonstrated the potential of ML algorithms to identify unusual patterns in IoT network traffic and provide adequate responses to emerging threats. For instance, the work [8] presented a comprehensive study on the use of ML for anomaly detection in IoT systems, emphasizing the ability to detect unauthorized access and other security breaches accurately. The authors of [9] expanded upon this by proposing a multi-layered machine learning approach for intrusion detection that significantly outperformed traditional rule-based systems, showing superior results in identifying complex attack patterns and reducing false positives. Beyond conventional ML techniques, deep learning models such as Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) have been extensively utilized for intrusion detection in IoT networks. For example, the authors of [10] proposed an autoencoder-based deep learning model that effectively identifies zero-day attacks while maintaining a low false-positive rate. Another significant work [11] integrated reinforcement learning with anomaly detection, allowing IoT security systems to adapt dynamically to new attack patterns without extensive retraining. In parallel, applying LLMs to cybersecurity has garnered increasing interest due to their ability to handle complex, unstructured data. Patel et al. introduced an LLM-based approach to network traffic analysis, utilizing prompt engineering to improve the model’s ability to identify malicious activities in IoT networks. Their findings suggested that LLMs could achieve results comparable to deep learning models, but with reduced computational overhead and improved resource efficiency [12]. Similarly, the authors of [13] applied LLMs for anomaly detection in IoT systems, achieving notable success in detecting known and novel threats. Their study demonstrated that LLMs could be trained to recognize the subtle differences in IoT traffic patterns, which are often missed by traditional models, thereby offering more robust security solutions. Another critical advancement in IoT security is the adoption of Federated Learning (FL) for distributed threat detection. Since IoT devices generate vast amounts of decentralized data, transmitting all traffic logs to a centralized server introduces privacy risks and scalability challenges. Federated learning addresses this by enabling devices to collaboratively train models without sharing raw data [14].

### v2

@@ Abstract
The increasing complexity and scale of the Internet of Things (IoT) have made security a critical concern. This paper presents a novel Large Language Model (LLM)-based framework for comprehensive threat detection and prevention in IoT environments. The system integrates lightweight LLMs fine-tuned on IoT-specific datasets (IoT-23, TON_IoT) for real-time anomaly detection and automated, context-aware mitigation strategies optimized for resource-constrained devices. A modular Docker-based deployment enables scalable and reproducible evaluation across diverse network conditions. Experimental results in simulated IoT environments demonstrate significant improvements in detection accuracy, response latency, and resource efficiency over traditional security methods. The proposed framework highlights the potential of LLM-driven, autonomous security solutions for future IoT ecosystems.
@@ I Introduction
The rapid growth of Internet of Things (IoT) ecosystems has introduced complex challenges in managing and optimizing system performance. IoT environments generate vast amounts of data, requiring efficient real-time processing and advanced operations such as resource allocation, predictive maintenance, and energy management. The diverse and dynamic nature of IoT devices further complicates these challenges, as devices often operate under heterogeneous conditions with varying computational and network capabilities [1]. The rapid expansion of IoT ecosystems introduces growing complexities in securing large-scale networks [2]. Traditional IoT security techniques often struggle to detect evolving and sophisticated threats. Recent advances in Large Learning Models (LLMs) have shown promise in improving security systems, offering adaptability to evolving threats, scalability across diverse IoT environments, and effectiveness in managing complex datasets. Recent studies demonstrate the growing potential of LLMs in cybersecurity applications. For example, the authors of [3] explored using domain-specific fine-tuning in LLMs for detecting zero-day vulnerabilities. Similarly, the authors of [4] proposed a hybrid approach combining LLMs with graph-based anomaly detection to secure IoT networks against emerging threats. The integration of LLMs with federated learning frameworks has also been successful in enhancing data privacy while maintaining robust security in distributed IoT environments as shown by the work [5]. Moreover, the authors in [6] highlighted the effectiveness of context-aware LLMs in identifying complex attack patterns specific to IoT ecosystems. The work [7] underlined the importance of adaptive learning techniques for LLMs to ensure continuous improvement in threat detection and prevention. Our model presents an LLM-based security system tailored to IoT environments, focusing on a two-component approach for threat detection and prevention. The proposed system integrates the strengths of LLMs, utilizing prompt engineering, fine-tuning, and domain-specific training to identify and respond to security breaches in real-time. By leveraging IoT-specific datasets such as IoT-23 and TON_IoT, the system continuously adapts to new threats, ensuring it remains effective in dynamic environments. This framework represents a significant step toward overcoming the limitations of traditional solutions while addressing the unique challenges of integrating advanced AI models into IoT ecosystems. The main contributions of this work are the following:
@@ II Related Works
In recent years, IoT security has been an active research area, with numerous studies exploring Machine Learning (ML) techniques for anomaly detection and intrusion prevention. Several works have demonstrated the potential of ML algorithms to identify unusual patterns in IoT network traffic and provide adequate responses to emerging threats. For instance, the work [8] presented a comprehensive study on the use of ML for anomaly detection in IoT systems, emphasizing the ability to detect unauthorized access and other security breaches accurately. The authors of [9] expanded upon this by proposing a multi-layered machine learning approach for intrusion detection that significantly outperformed traditional rule-based systems, showing superior results in identifying complex attack patterns and reducing false positives. Beyond conventional ML techniques, deep learning models such as Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) have been extensively utilized for intrusion detection in IoT networks. For example, the authors of [10] proposed an autoencoder-based deep learning model that effectively identifies zero-day attacks while maintaining a low false-positive rate. Another significant work [11] integrated reinforcement learning with anomaly detection, allowing IoT security systems to adapt dynamically to new attack patterns without extensive retraining. In parallel, applying LLMs to cybersecurity has garnered increasing interest due to their ability to handle complex, unstructured data. The authors of [12] introduced an LLM-based approach to network traffic analysis, utilizing prompt engineering to improve the model’s ability to identify malicious activities in IoT networks. Their findings suggested that LLMs could achieve results comparable to deep learning models, but with reduced computational overhead and improved resource efficiency. Another critical advancement in IoT security is the adoption of Federated Learning (FL) for distributed threat detection. Since IoT devices generate vast amounts of decentralized data, transmitting all traffic logs to a centralized server introduces privacy risks and scalability challenges. Federated learning addresses this by enabling devices to collaboratively train models without sharing raw data [13].